OpenStudy (anonymous):

Solve the Inequality: shome me the steps please ( -5)/(4-x^2) ≥0

OpenStudy (anonymous):

Show me**

OpenStudy (campbell_st):

look at the denominator solve 4 - x^2 = 0, x= 2 pr -2 values larger than 2, or -2 give a negative denominator... and with a negative numerator will result in a positive number... so x>= 2 or x<= -2

OpenStudy (anonymous):

\[-20+5x^2\ge0\] \[5x^2\le20\]\[x^2=20/5\] \[x^2=4\] \[x=\sqrt{4}\] \[x=\pm2\]